OpenStudy (anonymous):

Which of the following are geometric sequences? More than one answer will work. A. 2, -2, 2, -2, 2, -2, 2 B. 1, 4, 9, 16, 25, 36, 49 C. 10, 5, 2.5, 1.25, 0.625, 0.3125 D. 1, 2, 4, 8, 16, 32

OpenStudy (amistre64):

do you recall what defines a seq as geometric?

OpenStudy (anonymous):

connected by a common #

OpenStudy (amistre64):

given a seq: a,b,c,d a geometric seq is defined if: a*r = b b*r = c where r is a common ration, or multiplier c*r = d

OpenStudy (amistre64):

as such: d/c = c/b = b/a

OpenStudy (amistre64):

lets look at A A. 2, -2, 2, -2, 2, -2, 2 ; *-1 gives us each next value

OpenStudy (amistre64):

2/-2 = -2/2 = 2/-1 = ....

OpenStudy (anonymous):

I think C is one

OpenStudy (amistre64):

lol, that -1 just kinda slipped in there on accident

OpenStudy (amistre64):

yes, it looks like C is a /2 each time

OpenStudy (anonymous):

I dont think A is one

OpenStudy (amistre64):

why not?

OpenStudy (anonymous):

because the # just changes from 2 to -2 it just doesnt make sense to me

OpenStudy (amistre64):

2*r = -2 ; r=-1 -2*r = 2 ; r= -1 they have a common multiplier

OpenStudy (anonymous):

ohhh i see so A is one

OpenStudy (amistre64):

can you tell if D is one?

OpenStudy (anonymous):

I think it one because each time it gets multiplied by 2

OpenStudy (amistre64):

very good

OpenStudy (amistre64):

what about B? is there any hope for it?

OpenStudy (anonymous):

I dont think B is one because it doesnt look like it has any common #s

OpenStudy (anonymous):

32/16=16/8=8/4=4/2=2/1=2????????

OpenStudy (amistre64):

i agree 1*r = 4 ; r=4/1 4*r = 9 ; r= 9/4 since 4/1 not equal 9/4, there is no common value

OpenStudy (anonymous):

ok thank you so much! :)

OpenStudy (amistre64):

youre welcome